# DiCo: Revitalizing ConvNets for Scalable and Efficient Diffusion Modeling

### v1

@@ Abstract
Diffusion Transformer (DiT), a promising diffusion model for visual generation, demonstrates impressive performance but incurs significant computational overhead. Intriguingly, analysis of pre-trained DiT models reveals that global self-attention is often redundant, predominantly capturing local patterns—highlighting the potential for more efficient alternatives. In this paper, we revisit convolution as an alternative building block for constructing efficient and expressive diffusion models. However, naively replacing self-attention with convolution typically results in degraded performance. Our investigations attribute this performance gap to the higher channel redundancy in ConvNets compared to Transformers. To resolve this, we introduce a compact channel attention mechanism that promotes the activation of more diverse channels, thereby enhancing feature diversity. This leads to Diffusion ConvNet (DiCo), a family of diffusion models built entirely from standard ConvNet modules, offering strong generative performance with significant efficiency gains. On class-conditional ImageNet benchmarks, DiCo outperforms previous diffusion models in both image quality and generation speed. Notably, DiCo-XL achieves an FID of 2.05 at 256256 resolution and 2.53 at 512512, with a 2.7 and 3.1 speedup over DiT-XL/2, respectively. Furthermore, our largest model, DiCo-H, scaled to 1B parameters, reaches an FID of 1.90 on ImageNet 256256—without any additional supervision during training.
@@ 1 Introduction
Early diffusion models (e.g., ADM [11] and Stable Diffusion [61]) primarily employed hybrid U-Net [62] architectures that integrate convolutional layers with self-attention. More recently, Transformers [77] have emerged as a more powerful and scalable backbone [56, 3], prompting a shift toward fully Transformer-based designs. As a result, Diffusion Transformers (DiTs) are gradually supplanting traditional U-Nets, as seen in leading diffusion models such as Stable Diffusion 3 [16], FLUX [44], and Sora [6]. However, the quadratic computational complexity of self-attention presents substantial challenges, especially for high-resolution image synthesis. Recent efforts [90, 73, 19, 57, 84] have explored more efficient alternatives, focusing on linear-complexity RNN-like architectures, such as Mamba [20] and Gated Linear Attention [85]. While these models improve efficiency, their causal design inherently conflicts with the bidirectional nature of visual generation [23, 49], limiting their effectiveness. Furthermore, as illustrated in Fig. 3, even with highly optimized CUDA implementations, their runtime advantage over conventional DiTs remains modest in high-resolution settings. This leads us to a key question: Is it possible to design a hardware-efficient diffusion backbone that also preserves strong generative capabilities like DiTs?
@@
As shown in Fig. 2 and Fig. 3, DiCo models outperform state-of-the-art diffusion models on both the ImageNet 256256 and 512512 benchmarks. Notably, our DiCo-XL models achieve impressive FID scores of 2.05 and 2.53 at 256256 and 512512 resolution, respectively. In addition to performance gains, DiCo models exhibit considerable efficiency advantages over attention-based [77], Mamba-based [20], and linear attention-based [38] diffusion models. Specifically, at 256256 resolution, DiCo-XL achieves a 26.4% reduction in Gflops and is 2.7 faster than DiT-XL/2 [56]. At 512512 resolution, DiCo-XL operates 7.8 and 6.7 faster than the Mamba-based DiM-H [73] and DiS-H/2 [19] models, respectively. Our largest model, DiCo-H with 1 billion parameters, further reduces the FID on ImageNet 256256 to 1.90. These results collectively highlight the strong potential of DiCo in diffusion-based generative modeling.
@@
We conduct extensive experiments on the ImageNet 256256 and 512512 benchmarks. DiCo outperforms existing diffusion models in both generation quality and speed. To the best of our knowledge, this is the first work to demonstrate that well-designed fully convolutional backbones can achieve SOTA performance in diffusion-based generative modeling.
@@ 2 Related Work
Architecture of Diffusion Models. Early diffusion models commonly employ U-Net [62] as the foundational architecture [11, 28, 61]. More recently, a growing body of research has explored Vision Transformers (ViTs) [15] as alternative backbones for diffusion models, yielding remarkable results [56, 3, 52, 88, 59, 48]. Notably, DiT [56] has demonstrated the excellent performance of transformer-based architectures, achieving SOTA performance on ImageNet generation. However, the quadratic computational complexity inherent in ViTs presents significant challenges in terms of efficiency for long sequence modeling. To mitigate this, recent studies have explored the use of RNN-like architectures with linear complexity, such as Mamba [20] and linear attention [38], as backbones for diffusion models [19, 90, 73, 84, 57]. DiS [19] and DiM [73] employ Mamba to reduce computational overhead, while DiG [90] leverages Gated Linear Attention [85] to achieve competitive performance with improved efficiency. In this work, we revisit ConvNets as backbones for diffusion models. We show that, with proper design, pure convolutional architectures can achieve SOTA generative performance, providing an efficient and powerful alternative to DiTs.
@@ 4.1 Experimental Setup
| Model | Gflops | Throughput (it/s) | FID | IS | Precision | Recall |
| ADM-U [11] | 742 | - | 3.94 | 215.84 | 0.83 | 0.53 |
| LDM-4 [61] | - | - | 3.95 | 178.22 | 0.81 | 0.55 |
| U-ViT-H/2 [3] | 133.25 | 73.45 | 2.29 | 263.88 | 0.82 | 0.57 |
| Mamba-based diffusion models. | ||||||
| DiM-H [73] | 210 | 25.06 | 2.21 | - | - | - |
| DiS-H/2 [19] | - | 33.95 | 2.10 | 271.32 | 0.82 | 0.58 |
| DiffuSSM-XL [84] | 280.3 | - | 2.28 | 259.13 | 0.86 | 0.56 |
| DiMSUM-L/2 [57] | 84.49 | 59.13 | 2.11 | - | - | 0.59 |
| Baselines and Ours. | ||||||
| DiT-S/2 (400K) [56] | 6.06 | 1234.01 | 68.40 | - | - | - |
| DiC-S (400K) [75] | 5.9 | - | 58.68 | 25.82 | - | - |
| DiG-S/2 (400K) [90] | 4.30 | 961.24 | 62.06 | 22.81 | 0.39 | 0.56 |
| DiCo-S (400K) | 4.25 | 1695.73 | 49.97 | 31.38 | 0.48 | 0.58 |
| DiT-B/2 (400K) | 23.02 | 380.11 | 43.47 | - | - | - |
| DiC-B (400K) | 23.5 | - | 32.33 | 48.72 | - | - |
| DiG-B/2 (400K) | 17.07 | 345.89 | 39.50 | 37.21 | 0.51 | 0.63 |
| DiCo-B (400K) | 16.88 | 822.97 | 27.20 | 56.52 | 0.60 | 0.61 |
| DiT-L/2 (400K) | 80.73 | 114.63 | 23.33 | - | - | - |
| DiG-L/2 (400K) | 61.66 | 109.01 | 22.90 | 59.87 | 0.60 | 0.64 |
| DiCo-L (400K) | 60.24 | 288.32 | 13.66 | 91.37 | 0.69 | 0.61 |
| DiT-XL/2 (400K) | 118.66 | 76.90 | 19.47 | - | - | - |
| DiC-XL (400K) | 116.1 | - | 13.11 | 100.15 | - | - |
| DiG-XL/2 (400K) | 89.40 | 71.74 | 18.53 | 68.53 | 0.63 | 0.64 |
| DiCo-XL (400K) | 87.30 | 208.47 | 11.67 | 100.42 | 0.71 | 0.61 |
| DiT-XL/2 (w/ CFG) | 118.66 | 76.90 | 2.27 | 278.24 | 0.83 | 0.57 |
| DiG-XL/2 (w/ CFG) | 89.40 | 71.74 | 2.07 | 278.95 | 0.82 | 0.60 |
| DiCo-XL (w/ CFG) | 87.30 | 208.47 | 2.05 | 282.17 | 0.83 | 0.59 |
| DiC-H (w/ CFG) | 204.4 | - | 2.25 | - | - | - |
| DiCo-H (w/ CFG) | 194.15 | 117.57 | 1.90 | 284.31 | 0.83 | 0.61 |
| Model | Gflops | Throughput (it/s) | FID | IS | Precision | Recall |
| ADM-U [11] | 2813 | - | 3.85 | 221.72 | 0.84 | 0.53 |
| U-ViT-L/4 [3] | 76.52 | 128.49 | 4.67 | 213.28 | 0.87 | 0.45 |
| U-ViT-H/4 [3] | 133.27 | 73.42 | 4.05 | 263.79 | 0.84 | 0.48 |
| Mamba-based diffusion models. | ||||||
| DiM-H [73] | 708 | 7.39 | 3.78 | - | - | - |
| DiS-H/2 [19] | - | 8.59 | 2.88 | 272.33 | 0.84 | 0.56 |
| DiffuSSM-XL [84] | 1066.2 | - | 3.41 | 255.06 | 0.85 | 0.49 |
| Baselines and Ours. | ||||||
| DiT-XL/2 (1.3M) [56] | 524.70 | 18.58 | 13.78 | - | - | - |
| DiCo-XL (1.3M) | 349.78 | 57.45 | 8.10 | 132.85 | 0.78 | 0.62 |
| DiT-XL/2 (3M) | 524.70 | 18.58 | 12.03 | 105.25 | 0.75 | 0.64 |
| DiCo-XL (3M) | 349.78 | 57.45 | 7.48 | 146.35 | 0.78 | 0.63 |
| DiT-XL/2 (w/ CFG) | 524.70 | 18.58 | 3.04 | 240.82 | 0.84 | 0.54 |
| DiCo-XL (w/ CFG) | 349.78 | 57.45 | 2.53 | 275.74 | 0.83 | 0.56 |
@@ 4.2 Main Results
Comparison under the DiT Setting. In addition to DiT [56], we also select recent state-of-the-art diffusion models, DiG [90] and DiC [75], as baselines, since they similarly follow the experimental setup of DiT. Table 1 presents the comparison results on ImageNet 256256. Across different model scales trained for 400K iterations, our DiCo consistently achieves the best or second-best performance across all metrics. Furthermore, when using classifier-free guidance (CFG), our DiCo-XL achieves a state-of-the-art FID of 2.05 and an IS of 282.17. Beyond performance improvements, DiCo also demonstrates significant efficiency gains compared to both the baselines and Mamba-based models.
Table 2 presents the results on ImageNet 512512. At higher resolutions, our model demonstrates greater improvements in both performance and efficiency. Specifically, DiCo-XL achieves a state-of-the-art FID of 2.53 and an IS of 275.74, while reducing Gflops by 33.3% and achieving a 3.1 speedup compared to DiT-XL/2. These results highlight that our convolutional architecture remains highly efficient and effective for high-resolution image generation.
Scaling the model up. To further explore the potential of our model, we scale it up to 1 billion parameters. As shown in Table. 1, compared to DiCo-XL, the larger DiCo-H model achieves further improvements in FID (1.90 vs. 2.05), demonstrating the great scalability of our architecture. We provide additional scalability analysis and comparison results in Appendix Sec. C and Sec. D. More generated samples can be found in Appendix Sec. E and the submitted Supplementary Material.
@@ 4.3 Ablation Study
For the ablation study, we use the small-scale model and evaluate performance on the ImageNet 256256 benchmark. All models are trained for 400K iterations. In this section, self-attention in DiT is not accelerated using FlashAttention-2 to ensure a fair comparison with other efficient attention mechanisms. We analyze both the overall architecture and the contributions of individual components within DiCo to better understand their impact on model performance.
@@
Component Ablation. We conduct a component-wise analysis of DiCo, examining the effects of the activation function, convolutional kernel size, compact channel attention (CCA), and the conv module (CM). The overall ablation results are summarized in Table 4. Increasing the convolutional kernel size leads to further performance gains but at the expense of reduced efficiency, highlighting a trade-off between performance and computational cost. The introduction of CCA results in a 4.81-point improvement in FID. As illustrated by the feature visualizations in Fig. 7, CCA significantly enhances feature diversity, demonstrating its effectiveness in improving the model’s representational capacity. We also compare CCA with SE module [34] and Channel-wise Self-Attention [89]; despite its simplicity, CCA achieves superior performance and higher efficiency. For the Conv Module, we benchmark it against several advanced efficient attention mechanisms (Window Attention [50], Focused Linear Attention [22], Agent Attention [59]). The results show that our CM offers both better performance and computational efficiency.
| Model | Skip | Hierarchy | #Params | Gflops | Throughput | FID | IS | Pre | Rec |
@@
| Model | #Params | Gflops | Throughput | FID | IS | Pre | Rec |

### v2

@@ Abstract
Diffusion Transformer (DiT), a promising diffusion model for visual generation, demonstrates impressive performance but incurs significant computational overhead. Intriguingly, analysis of pre-trained DiT models reveals that global self-attention is often redundant, predominantly capturing local patterns—highlighting the potential for more efficient alternatives. In this paper, we revisit convolution as an alternative building block for constructing efficient and expressive diffusion models. However, naively replacing self-attention with convolution typically results in degraded performance. Our investigations attribute this performance gap to the higher channel redundancy in ConvNets compared to Transformers. To resolve this, we introduce a compact channel attention mechanism that promotes the activation of more diverse channels, thereby enhancing feature diversity. This leads to Diffusion ConvNet (DiCo), a family of diffusion models built entirely from standard ConvNet modules, offering strong generative performance with significant efficiency gains. On class-conditional ImageNet generation benchmarks, DiCo-XL achieves an FID of 2.05 at 256256 resolution and 2.53 at 512512, with a 2.7 and 3.1 speedup over DiT-XL/2, respectively. Furthermore, experimental results on MS-COCO demonstrate that the purely convolutional DiCo exhibits strong potential for text-to-image generation.
@@ 1 Introduction
Early diffusion models (e.g., ADM [14] and Stable Diffusion [67]) primarily employed hybrid U-Net [68] architectures that integrate convolutional layers with self-attention. More recently, Transformers [83] have emerged as a more powerful and scalable backbone [62, 6], prompting a shift toward fully Transformer-based designs. As a result, Diffusion Transformers (DiTs) are gradually supplanting traditional U-Nets, as seen in leading diffusion models such as Stable Diffusion 3 [20], FLUX [49], and Sora [9]. However, the quadratic computational complexity of self-attention presents substantial challenges, especially for high-resolution image synthesis. Recent efforts [100, 79, 25, 63, 91, 2] have explored more efficient alternatives, focusing on linear-complexity RNN-like architectures, such as Mamba [26] and Gated Linear Attention [92]. While these models improve efficiency, their causal design inherently conflicts with the bidirectional nature of visual generation [30, 55], limiting their effectiveness. Furthermore, as illustrated in Fig. 3, even with highly optimized CUDA implementations, their runtime advantage over conventional DiTs remains modest in high-resolution settings. This leads us to a key question: Is it possible to design a hardware-efficient diffusion backbone that also preserves strong generative capabilities like DiTs?
@@
As shown in Fig. 2 and Fig. 3, DiCo models outperform recent diffusion models on both the ImageNet 256256 and 512512 benchmarks. Notably, our DiCo-XL models achieve impressive FID scores of 2.05 and 2.53 at 256256 and 512512 resolution, respectively. In addition to performance gains, DiCo models exhibit considerable efficiency advantages over attention-based [83], Mamba-based [26], and linear attention-based [44] diffusion models. Specifically, at 256256 resolution, DiCo-XL achieves a 26.4% reduction in Gflops and is 2.7 faster than DiT-XL/2 [62]. At 512512 resolution, DiCo-XL operates 7.8 and 6.7 faster than the Mamba-based DiM-H [79] and DiS-H/2 [25] models, respectively. Our largest model, DiCo-H with 1 billion parameters, further reduces the FID on ImageNet 256256 to 1.90. In addition, we validate the applicability of DiCo for text-to-image generation on the MS-COCO dataset. These results collectively highlight the strong potential of DiCo in diffusion-based generative modeling.
@@
We conduct extensive experiments on class-conditional ImageNet benchmarks. DiCo outperforms recent diffusion models in both generation quality and speed. Furthermore, the purely convolutional DiCo demonstrates strong potential in text-to-image generation.
@@ 2 Related Work
Architecture of Diffusion Models. Early diffusion models commonly employ U-Net [68] as the foundational architecture [14, 34, 67]. More recently, a growing body of research has explored Vision Transformers (ViTs) [18] as alternative backbones for diffusion models, yielding remarkable results [62, 6, 58, 96, 65, 54]. Notably, DiT [62] has demonstrated the excellent performance of transformer-based architectures, achieving SOTA performance on ImageNet generation. However, the quadratic computational complexity inherent in ViTs presents significant challenges in terms of efficiency for long sequence modeling. To mitigate this, recent studies have explored the use of RNN-like architectures with linear complexity, such as Mamba [26] and linear attention [44], as backbones for diffusion models [25, 100, 79, 91, 63]. DiS [25] and DiM [79] employ Mamba to reduce computational overhead, while DiG [100] leverages Gated Linear Attention [92] to achieve competitive performance with improved efficiency. In this work, we revisit ConvNets as backbones for diffusion models. We show that, with proper design, pure convolutional architectures can achieve superior generative performance, providing an efficient and powerful alternative to DiTs.
@@ 3.3 DiCo Block
Modification for text-to-image. We investigate two different approaches for incorporating textual features into DiCo. The first uses the widely adopted cross-attention [12] mechanism, integrated into the DiCo architecture to fuse text and visual features. The second transforms CLIP text embeddings into dynamic depthwise convolution (DWC) kernels. We pad the text embeddings to a length of 81, feed them through a learnable MLP, and reshape the output into a kernel. This kernel dynamically modulates DiCo’s features via depthwise convolution. In this way, we can construct a fully convolutional text-to-image DiCo model without relying on any self-attention or cross-attention operations. We provide its detailed PyTorch implementation in Algorithm 1. Both feature fusion modules are inserted after the Conv Module within each DiCo block.
| Model | Token Mixing Type | Gflops | Throughput | FID | IS | Pre. | Rec. |
| ADM-U [14] | Conv + Attn | 742 | - | 3.94 | 215.84 | 0.83 | 0.53 |
| LDM-4 [67] | Conv + Attn | - | - | 3.95 | 178.22 | 0.81 | 0.55 |
| U-ViT-H/2 [6] | Attn | 133.25 | 73.45 | 2.29 | 263.88 | 0.82 | 0.57 |
| Mamba-based diffusion models. | |||||||
| DiM-H [79] | Conv + SSM | 210 | 25.06 | 2.21 | - | - | - |
| DiS-H/2 [25] | Conv + SSM | - | 33.95 | 2.10 | 271.32 | 0.82 | 0.58 |
| DiffuSSM-XL [91] | SSM | 280.3 | - | 2.28 | 259.13 | 0.86 | 0.56 |
| DiMSUM-L/2 [63] | Attn + SSM | 84.49 | 59.13 | 2.11 | - | - | 0.59 |
| Baselines and Ours (w/ the same hyperparameters). | |||||||
| DiT-S/2 (400K) [62] | Attn | 6.06 | 1234.01 | 68.40 | - | - | - |
| DiC-S (400K) [81] | Conv | 5.9 | - | 58.68 | 25.82 | - | - |
| DiG-S/2 (400K) [100] | Conv + Attn | 4.30 | 961.24 | 62.06 | 22.81 | 0.39 | 0.56 |
| DiCo-S (400K) | Conv | 4.25 | 1695.73 | 49.97 | 31.38 | 0.48 | 0.58 |
| DiT-B/2 (400K) | Attn | 23.02 | 380.11 | 43.47 | - | - | - |
| DiC-B (400K) | Conv | 23.5 | - | 32.33 | 48.72 | - | - |
| DiG-B/2 (400K) | Conv + Attn | 17.07 | 345.89 | 39.50 | 37.21 | 0.51 | 0.63 |
| DiCo-B (400K) | Conv | 16.88 | 822.97 | 27.20 | 56.52 | 0.60 | 0.61 |
| DiT-L/2 (400K) | Attn | 80.73 | 114.63 | 23.33 | - | - | - |
| DiG-L/2 (400K) | Conv + Attn | 61.66 | 109.01 | 22.90 | 59.87 | 0.60 | 0.64 |
| DiCo-L (400K) | Conv | 60.24 | 288.32 | 13.66 | 91.37 | 0.69 | 0.61 |
| DiT-XL/2 (400K) | Attn | 118.66 | 76.90 | 19.47 | - | - | - |
| DiC-XL (400K) | Conv | 116.1 | - | 13.11 | 100.15 | - | - |
| DiG-XL/2 (400K) | Conv + Attn | 89.40 | 71.74 | 18.53 | 68.53 | 0.63 | 0.64 |
| DiCo-XL (400K) | Conv | 87.30 | 208.47 | 11.67 | 100.42 | 0.71 | 0.61 |
| DiT-XL/2 (w/ CFG) | Attn | 118.66 | 76.90 | 2.27 | 278.24 | 0.83 | 0.57 |
| DiG-XL/2 (w/ CFG) | Conv + Attn | 89.40 | 71.74 | 2.07 | 278.95 | 0.82 | 0.60 |
| DiCo-XL (w/ CFG) | Conv | 87.30 | 208.47 | 2.05 | 282.17 | 0.83 | 0.59 |
| DiC-H (w/ CFG) | Conv | 204.4 | - | 2.25 | - | - | - |
| DiCo-H (w/ CFG) | Conv | 194.15 | 117.57 | 1.90 | 284.31 | 0.83 | 0.61 |
@@ 4.1 Experimental Setup
| Model | Token Mixing Type | Gflops | Throughput | FID | IS | Pre. | Rec. |
| ADM-U [14] | Conv + Attn | 2813 | - | 3.85 | 221.72 | 0.84 | 0.53 |
| U-ViT-L/4 [6] | Attn | 76.52 | 128.49 | 4.67 | 213.28 | 0.87 | 0.45 |
| U-ViT-H/4 [6] | Attn | 133.27 | 73.42 | 4.05 | 263.79 | 0.84 | 0.48 |
| Mamba-based diffusion models. | |||||||
| DiM-H [79] | Conv + SSM | 708 | 7.39 | 3.78 | - | - | - |
| DiS-H/2 [25] | Conv + SSM | - | 8.59 | 2.88 | 272.33 | 0.84 | 0.56 |
| DiffuSSM-XL [91] | Attn + SSM | 1066.2 | - | 3.41 | 255.06 | 0.85 | 0.49 |
| Baselines and Ours (w/ the same hyperparameters). | |||||||
| DiT-XL/2 (1.3M) [62] | Attn | 524.70 | 18.58 | 13.78 | - | - | - |
| DiCo-XL (1.3M) | Conv | 349.78 | 57.45 | 8.10 | 132.85 | 0.78 | 0.62 |
| DiT-XL/2 (3M) | Attn | 524.70 | 18.58 | 12.03 | 105.25 | 0.75 | 0.64 |
| DiCo-XL (3M) | Conv | 349.78 | 57.45 | 7.48 | 146.35 | 0.78 | 0.63 |
| DiT-XL/2 (w/ CFG) | Attn | 524.70 | 18.58 | 3.04 | 240.82 | 0.84 | 0.54 |
| DiCo-XL (w/ CFG) | Conv | 349.78 | 57.45 | 2.53 | 275.74 | 0.83 | 0.56 |
@@ 4.2 Main Results
Comparison under the DiT Setting. In addition to DiT [62], we also select recent diffusion models, DiG [100] and DiC [81], as baselines, since they similarly follow the experimental setup of DiT.Table 1 presents the comparison results on ImageNet 256256. Across different model scales trained for 400K iterations, our DiCo consistently achieves the best or second-best performance across all metrics. Furthermore, when using classifier-free guidance (CFG), our DiCo-XL achieves an FID of 2.05 and an IS of 282.17. Beyond performance improvements, DiCo also demonstrates significant efficiency gains compared to both the baselines and Mamba-based models.
Table 2 presents the results on ImageNet 512512. At higher resolutions, our model demonstrates greater improvements in both performance and efficiency. Specifically, DiCo-XL achieves an FID of 2.53 and an IS of 275.74, while reducing Gflops by 33.3% and achieving a 3.1 speedup compared to DiT-XL/2. These results highlight that our convolutional architecture remains highly efficient and effective for high-resolution image generation.
Scaling the model up. To further explore the potential of our model, we scale it up to 1 billion parameters. As shown in Table 1, compared to DiCo-XL, the larger DiCo-H achieves further improvements in FID (1.90 vs. 2.05), demonstrating the great scalability of our architecture. More comparison results and generated images can be found in the Appendix Sec. D and Sec. E.
Text-to-Image Generation. We follow [6] to conduct small-scale text-to-image generation experiments. Specifically, we adopt the same experimental setup as [96]: training and evaluating models from scratch on MS-COCO [53], using CLIP as the text encoder with a token length of 77.
As shown in Table 3, our DiCo achieves superior generation quality for text-to-image generation. Notably, using text conditional DWC in place of cross-attention further improves throughput while maintaining competitive performance. This suggests that the fully convolutional DiCo has the potential to serve as the backbone for large-scale text-to-image diffusion models.
@@ 4.3 Ablation Study
For the ablation study, we use the small-scale model and evaluate performance on the ImageNet 256256 benchmark to enable fast training speed. All models are trained for 400K iterations and evaluated without CFG. Notably, in this section, self-attention in DiT is not accelerated using FlashAttention-2 to ensure a fair speed comparison with other efficient attention mechanisms.
| Model | Type | FID |
| AttnGAN [90] | GAN | 35.49 |
| DM-GAN [101] | GAN | 32.64 |
| VQ-Diffusion [27] | Discrete Diffusion | 19.75 |
| DF-GAN [78] | GAN | 19.32 |
| XMC-GAN [98] | GAN | 9.33 |
| Frido [24] | Diffusion | 8.97 |
| LAFITE [99] | GAN | 8.12 |
| U-Net [6] | Diffusion | 7.32 |
| U-ViT-S/2 [6] | Diffusion | 5.95 |
| U-ViT-S/2 (Deep) [6] | Diffusion | 5.48 |
| MMDiT [96] | Diffusion | 5.30 |
| MMDiT+REPA [96] | Diffusion | 4.14 |
| DiCo-CrossAttn | Diffusion | 4.87 |
| DiCo-DWC | Diffusion | 4.93 |
We analyze both the overall architecture and the contributions of individual components within DiCo to better understand their impact on model performance.
@@
Component-wise Ablation. We conduct a component-wise analysis of DiCo, examining the effects of the activation function, convolutional kernel size, compact channel attention (CCA), and the conv module (CM). The overall ablation results are summarized in Table 5. Increasing the convolutional kernel size leads to further performance gains but at the expense of reduced efficiency, highlighting a trade-off between performance and computational cost.
The introduction of CCA results in a 4.81-point improvement in FID. As illustrated in Fig. 7, CCA significantly enhances feature diversity, demonstrating its effectiveness in improving the model’s representational capacity. We also compare CCA with SE module [40] and Channel-wise Self-Attention [97]; despite its simplicity, CCA achieves superior performance and higher efficiency.
For the Conv Module, we benchmark it against several advanced efficient attention mechanisms (Window Attention [56], Focused Linear Attention [29], Agent Attention [65]). The results show that our CM offers both better performance and computational efficiency.
| Model | Skip | Hierarchy | #Params | Gflops | Throughput | FID | IS | Pre. | Rec. |
@@
| Model | #Params | Gflops | Throughput | FID | IS | Pre. | Rec. |